OpenStudy (kkrpbra):

A team of horses travels 1,680 miles from California to Missouri to deliver a letter. The team completes the trip in 10 days. What was the team's average speed during the trip? Include any rest stops in the calculation. One day is the same as 24 hours.

OpenStudy (anonymous):

10 days * 24 hours = 240 hours. 1680 miles / 240 hours = 168/24 = 84/12 = 42/6 = 21/3 = 7 miles/hour, or 7mph.

OpenStudy (anonymous):

How do we include reststops ?

OpenStudy (anonymous):

Don't worry about the rest stops, that's a distraction from the real question.

OpenStudy (anonymous):

By which I mean, they're already included: You know how long the team took to get there, and how far they traveled.